# Advancing Language Multi-Agent Learning with Credit Re-Assignment for Interactive Environment Generalization

### v2

@@ Abstract
11footnotetext: Equal contribuation.22footnotetext: Work done at Institute for AI Industry Research (AIR), Tsinghua University.
@@ 1 Introduction
However, current efforts in language agent learning still face challenges to balance both performance and generalizability across interactive environments. (1) Single-agent learning methods (Chen et al., 2023; Gur et al., 2024; Furuta et al., 2024; Bai et al., 2024) heavily relies on in-domain supervised data or rewarding signals to improve environment-specific performance, which restricts its generalization across environments, such as transitioning between web environments using HTML and mobile environments using Android automator. (2) Despite being trained on vast amounts of data from diverse domains, single agents based on open-source LLMs (Zeng et al., 2024; Zhang et al., 2024a; Yin et al., 2024; Liu et al., 2025) demonstrate only moderate generalization capabilities and lag behind closed-source models. (3) Although multi-agent learning methods (Qiao et al., 2024; Liang et al., 2024) have better performance, existing ones are often constrained by rigid role assignments and lack dedicated designs for generalization, which limits their adaptability to unseen environments, e.g., an agent designed to retrieve documents for question answering is not feasible to handle mobile operations. In short, it is still unclear how to achieve strong performance and good generalization in interactive environments at the same time.
@@
In this work, we introduce a reinforcement learning framework for language multi-agent systems, named CollabUIAgents, designed to address challenges above in real-world interactive environments. The methodology is inspired from the success of classic multi-agent reinforcement learning (MARL) to stimulate collaboration (Ma & Luo, 2022). Beyond previous work (Tran et al., 2025), collaboration among langugae agents may also be beneficial for generalization across environments, as illustrated in Figure 1. Compared to existing methods using LLMs in credit assignment for non-language multi-agent settings (Qu et al., 2025; Lin et al., 2025), further considerations are made in our framework to enrich sparse rewarding signals, enhance generalization, and mitigate errors in credit assignment. We propose a novel multi-agent credit re-assignment (CR) strategy, assigning process rewards without using environment-specific outcome rewards, but with the world knowledge embedded in LLMs, and learning with synthesized preference data, aiming to foster generalizable, collaborative behaviors among the role-free agents’ policies.
Specially, an agentic fine-tuned model, as the base UIAgent, is used to initialize a multi-agent system, where each agent has its own policy. After rolling out actions from the policy multi-agent system, the critic agent allocates process rewards at both the agent and conversation round levels based on its comprehension of the environment state and agent behaviors. This approach not only enables finer-grained rewards without training numerous value estimators for agents, but also expands data scales by restoring generalizable behaviors from failed trajectories according to the training environment. To avoid misleading the agents with incorrect CR, policies are optimized with preference data synthesized by the adversarial agent, to ensure guiding them with correct preference signals. After preference optimization  (Rafailov et al., 2023), the multi-agent system is updated in both model parameters and edges of the communication structure. Empirically, we show that with the CR strategy, (1) preference optimization benefits performance and generalization; (2) edge updates is crucial to orchestrate trained LLMs in multi-agent systems.
@@
  • Extensive experiments show that CollabUIAgents surpasses the performance of strong baselines and shows competitiveness comparable to the guidance LLM of CR in both trained and target interactive environments, even under cross-environment generalization.
@@ 2 Formulation and Notations
Outcome Reward Function  The outcome reward $R_{o}\in\{0,1\}$ is defined in the environment based on static rules. Let the terminal step be $t^{*}$,
@@
Thus the outcome reward is sparse, as only the terminal state $s^{*}$ gives out positive rewards, posing a challenge to traditional RL approaches.
@@ 3 CollabUIAgents Framework
The CollabUIAgents framework is designed to achieve both high performance and generalizability in interactive environments. It optimizes language multi-agent systems without predefined roles. As shown in Figure 2, the base model undergoes agentic fine-tuning (detailed in Appendix B) and then MARL process. This section elaborates the multi-agent system architecture and the multi-agent learning process with multi-agent credit re-assignment.
@@ 3.1 Multi-Agent System Architecture
The architecture of the multi-agent system ($\mathcal{G}$) in CollabUIAgents is in consistency with previous works (Zhuge et al., 2024a; Liu et al., 2024), which consists of $|\mathcal{G}|=n$ agents, each represented by a policy $\pi_{i}$ that communicate with each other through a message network $\mathcal{E}_{\mathcal{G}}$. As shown in Figure 2, the network is a directed acyclic graph (DAG), where messages are passed from $\pi_{i_{1}}$ to $\pi_{i_{2}}$ if there is an edge pointing from $\pi_{i_{1}}$ to $\pi_{i_{2}}$. Specifically, the message is from the output of $\pi_{i_{1}}$. It is worth noting that the framework remains the compatibility for more complex architecture, which is left for future work. We instead use DAGs for simplicity.
@@ 3.2 Language Multi-Agent RL
As shown in Figure 2, the essential stage of CollabUIAgents is MARL, after agentic fine-tuning the base model into a base UIAgent (Appendix B). We introduce a novel multi-agent Credit Re-Assignment (CR) strategy for better MARL. This approach utilizes the critic agent to provide fine-grained rewards at both the agent and conversation round levels, enabling agents to learn effectively from general knowledge of LLMs rather than the sparse environment-specfic reward, and improve collaboration by preference optimization with synthesized data from the adversarial agent. We also introduce an Edge Update trick to better accommodate trained LLMs in the multi-agent system.
Credit Re-Assignment  Traditional MARL systems use value decomposition for credit assignment (Ma & Luo, 2022), and recent work also uses multiple value estimators to alleviate reward sparcity in interactive environments (Bai et al., 2024). However, training critic models for each agent based on outcome rewards is computationally expensive and may hinder generalization. Instead, by using an LLM-based critic agent, represented by $\pi_{\mathrm{critic}}$, we provide process rewards in a finer granularity for each agent at each conversation round. The action matrix $\bm{A}_{t}$ is assessed by the critic agent based on observed environment state $o_{t}$, interaction history $H_{t-1}$, and the task query $q$, generating a reward matrix $\bm{R}_{t}=(r_{t}^{i,j})$, $i=1,...,n,j=1,...,m$, where $r_{t}^{i,j}\in\{0,1\}$ denotes the intermediate reward for $a_{t}^{i,j}$:
@@
where $\pi_{\mathrm{critic}}(\cdot,\cdot,\cdot,\cdot)\in[0,1]^{n\times m}$. To overcome the potential incorrect CR results, we introduce an adversarial agent $\pi_{\mathrm{adv}}$ to synthesize preference data for policy optimization substituting value-based learning. Specially, $\pi_{\mathrm{adv}}$ generates low-quality responses $a_{t}^{i,j,-}$ paired with $a_{t}^{i,j}$:
@@
Here is the rationale of the design: (1) The critic agent provides a more detailed reward signal for each agent without training multiple value estimators; (2) The critic agent’s assessment is based on general pretrained LLMs, which by practice could expand the data scale by restoring intermediate decisions from failed trajectories according to the training environment, and thus may enhance performance; (3) Although errors in CR are inevitable, the synthesized preference data can still provide agents with meaningful preference signals. For example, even if $a_{t}^{i,j}$ is not the optimal action, it is still better than $a_{t}^{i,j,-}$. Indeed, environment-specific actions beyond general knowledge of the LLM might be misjudged. We argue that the situation is similar when a value estimator is poorly learned, where the policy might be optimized to a wrong direction, but our approach could keep rolling out new trajectories based on these actions without unlearning them. The qualitative study shown in Appendix A demonstrates CR results could be valid without the outcome reward.
MARL with Edge Updates  Different from classic MARL settings, agents in CollabUIAgents could communicate and the message network should also be rolled out in the optimization. To alleviate the overhead of learning the optimal network from enormous combinations of edges, we introduce an edge update trick, that randomly update edges to form a DAG message network, independently from model parameter updates. Through this process, we encourage agents to learn the awareness of multi-agent collaboration and adapt to diverse message networks rather than being rigid in locally optimal DAG pattern. As shown in Figure 2, the edge update is functioned before rolling out actions from the policy system. The overall learning objective for each agent $\pi_{i}$ is formulated as:
@@
where $\hat{H}_{t}^{i}=\{H_{t-1},C_{t}^{i,j},q\}$, $\theta_{i}$ are the updating parameters of agent $\pi_{i}$ and $\mathrm{ref}_{i}$ is the original model used as reference policy, $K_{|\mathcal{G}|}$ is a fully connected graph of $|\mathcal{G}|$ nodes, $\mathcal{E}_{\mathcal{G}}^{\prime}$ represents a DAG subgraph sampled from $K_{|\mathcal{G}|}$, and $\mathcal{P}(\mathcal{G},\mathcal{E}_{\mathcal{G}}^{\prime})$ is the preference dataset sampled with agents in the message network $\mathcal{E}_{\mathcal{G}}^{\prime}$. This objective encourages the policy $\pi_{i}$ to assign higher probabilities to preferred actions $a_{t}^{i,j}$ compared to adversarial actions $a_{t}^{i,j,-}$. The policy system is updated online and off-policy, similar to previous work (Bai et al., 2024).
@@ 3.3 Cross-Environment Adaptation
Direct Transfer  In scenarios where the new environment shares similarities with the training environment, agents can be directly deployed without additional training. For example, agents trained in mobile UI environments can directly apply their knowledge to web environments, leveraging the knowledge common interaction patterns and UI elements. The multi-agent setup may also decrease error rates through collaborations as expected.
Continual MARL  When the new environment presents significant differences or the higher success rates are demanded, agents can undergo further training using the MARL with the CR strategy on the new environment. This continual learning approach allows agents to refine their policies without stashing the knowledge of previous environments.
Method Base model #Params #Agents Input SRAndroidWorld SRMMiniWoB++ $\boldsymbol{\Delta}$Generalization
@@
M3A GPT-4 N/A 1 Text 30.6 59.7 -
M3A GPT-4 N/A 1 Text & Image 25.4 67.7 -
SeeAct GPT-4 N/A 1 Text & Image 15.5 66.1 -
M3A Gemini 1.5 Pro N/A 1 Text 19.4 57.4 -
M3A Gemini 1.5 Pro N/A 1 Text & Image 22.8 40.3 -
@@
Base Model Qwen2 7B 1 Text 06.2 12.9 -
Base Model Qwen2 VL 2B 1 Text & Image 00.0 10.0 -
InfiGUIAgent Qwen2 VL 2B 1 Text & Image 09.1 15.6 05.6
Base Model Qwen2.5 VL 3B 1 Text & Image 10.1 18.7 -
DigiRL Qwen2.5 VL 3B 1 Text & Image 22.3 35.2 16.5
DigiRL Qwen2.5 VL 3B 4 Text & Image 20.1 38.7 20.0
@@
Base UIAgent Qwen2 7B 1 Text 18.9 48.4 35.5
Base UIAgent Qwen2 7B 4 Text 21.4 53.2 40.3
CollabUIAgentsmobile Qwen2 7B 4 Text 29.3 61.2 48.3
Table 1: Experimental results on mobile operation environments. Success rates (SR) in AndoridWorld and MobileMiniWoB++ (MMiniWoB++) are listed. $\boldsymbol{\Delta}$Generalization indicates the performance gap between the base model and agent learning methods based on the model in MobileMiniWoB++.
@@ 4.1 Experimental Settings
Environments  We conduct experiments in both mobile operation and web browsing environments. For the mobile environments, we use AndroidWorld (Rawles et al., 2025) for training and MobileMiniWoB++ (Rawles et al., 2025) for testing: (1) AndroidWorld has 116 programmatic tasks across 20 real-world apps, such as Chrome, Markor, and Pro Expense. (2) MobileMiniWoB++ is derived from MiniWoB++ (Shi et al., 2017), which is a web-based benchmark. MobileMiniWoB++ shares the same observation space as AndroidWorld and supports 92 tasks from MiniWoB++. We use the success rate (SR) as an evaluation metric. For the web environments, we leverage Mind2Web (Deng et al., 2023) for training and AutoWebBench (Lai et al., 2024) for testing: (1) Mind2Web features over 2,000 open-ended tasks sourced from 137 websites in 31 different domains. (2) AutoWebBench is a bilingual benchmark featuring approximately 10,000 traces, from mainstream Chinese and English websites, providing a diverse dataset for web browsing. We use the step-success rate (SSR) as the evaluation metric. For agent learning methods on open-source LLMs, we use the performance gap between the base model and the trained model in unseen environments, $\boldsymbol{\Delta}$Generalization, to indicate generalizability.
@@
Method Base model #Params #Agents Input SSRMind2Web SSRAutoWebBench $\boldsymbol{\Delta}$Generalization

### v3

@@ Abstract
11footnotetext: Equal contribuation. Corresponding authors: Peng Li, Yang Liu.22footnotetext: Work done at Institute for AI Industry Research (AIR), Tsinghua University.
@@ 1 Introduction
However, current efforts in language agent learning still face challenges to balance both performance and generalizability across interactive environments. (1) Single-agent learning methods (Chen et al., 2023; Gur et al., 2024; Furuta et al., 2024; Bai et al., 2024) heavily relies on in-domain supervised data or rewarding signals to improve environment-specific performance, which restricts its generalization across environments, such as transitioning between web environments using HTML and mobile environments using Android automator. (2) Despite being trained on vast amounts of data from diverse domains, single agents based on open-source LLMs (Zeng et al., 2024; Zhang et al., 2024a; Yin et al., 2024; Liu et al., 2025a) demonstrate only moderate generalization capabilities and lag behind closed-source models. (3) Although multi-agent learning methods (Qiao et al., 2024; Liang et al., 2024) have better performance, existing ones are often constrained by rigid role assignments and lack dedicated designs for generalization, which limits their adaptability to unseen environments, e.g., an agent designed to retrieve documents for question answering is not feasible to handle mobile operations. In short, it is still unclear how to achieve strong performance and good generalization in interactive environments at the same time.
@@
In this work, we introduce a reinforcement learning framework for language multi-agent systems, named CollabUIAgents, designed to address the challenges above in real-world interactive environments. The methodology is inspired by the success of classic multi-agent reinforcement learning (MARL) to stimulate collaboration (Ma & Luo, 2022). Beyond previous work (Tran et al., 2025), collaboration among langugae agents may also be beneficial for generalization across environments, as illustrated in Figure 1. Compared to existing methods using LLMs in credit assignment for non-language multi-agent settings (Qu et al., 2025; Lin et al., 2025) (Appendix A), further considerations are made to enrich sparse outcome rewards, enhance adaptation across language environments, and enable multi-agent training for language agents. We propose a novel multi-agent credit re-assignment (CR) strategy, assigning process rewards without using environment-specific outcome rewards, but with the world knowledge embedded in LLMs, and learning with synthesized preference data, aiming to foster generalizable, collaborative behaviors.
The core of the framework is to be completely powered by MARL-driven policy learning, rather than fixed role prompts. Specially, an agentic fine-tuned model, as the base UIAgent, is used to initialize a multi-agent system, where each agent has its own policy. After rolling out actions from the policy multi-agent system, the critic agent allocates process rewards at both the agent and conversation round levels based on its comprehension of the environment state and agent behaviors. This approach not only enables finer-grained rewards without training numerous value estimators for agents, but also expands data scales by restoring generalizable behaviors from failed trajectories according to the training environment. To avoid misleading the agents with incorrect CR, policies are optimized with preference data synthesized by the adversarial agent, to ensure guiding them with correct preference signals. After preference optimization  (Rafailov et al., 2023), the multi-agent system is updated in both model parameters and edges of the communication structure. Empirically, we show that with the CR strategy, (1) preference optimization benefits performance and generalization; (2) edge updates is crucial to orchestrate trained LLMs in multi-agent systems.
@@
  • Extensive experiments show that CollabUIAgents surpasses the performance of strong baselines and shows competitiveness comparable to the guidance LLM of CR in both trained and target environments, even under cross-environment generalization.
@@ 2 Formulation and Notations
Outcome Reward Function  The outcome reward $R_{o}\in\{0,1\}$ is defined in the environment based on static rules. Static rules are predefined to check whether agents arrive in a successful terminal state inherent to the given task query. Specifically, in the following experiments, AndriodWorld and MobileMiniWoB++ (Rawles et al., 2025) feature online environments and according state annotations, and Mind2Web (Deng et al., 2023) and AutoWebBench (Lai et al., 2024) use offline trajectories and verifiers. Let the terminal step be $t^{*}$,
@@
Thus, the outcome reward is sparse, as only the terminal state $s^{*}$ gives out positive rewards, posing a challenge to traditional RL approaches.
@@ 3 CollabUIAgents Framework
The CollabUIAgents framework is designed to achieve both high performance and generalizability in interactive environments. It optimizes language multi-agent systems without predefined roles. As shown in Figure 2, the base model undergoes agentic fine-tuning (detailed in Appendix C) and then the MARL process. This section elaborates the multi-agent system architecture and the multi-agent learning process with multi-agent credit re-assignment.
@@ 3.1 Multi-Agent System Architecture
The architecture of the multi-agent system ($\mathcal{G}$) in CollabUIAgents is consistent with previous work (Zhuge et al., 2024a; Liu et al., 2024), which consists of $|\mathcal{G}|=n$ agents, each represented by a policy $\pi_{i}$ that communicate with each other through a message network $\mathcal{E}_{\mathcal{G}}$. As shown in Figure 2, the network is a directed acyclic graph (DAG), where messages are passed from $\pi_{i_{1}}$ to $\pi_{i_{2}}$ if there is an edge pointing from $\pi_{i_{1}}$ to $\pi_{i_{2}}$. Specifically, the message comes from the output of $\pi_{i_{1}}$. The framework remains the compatibility for more complex architectures, which is left for future work. We instead use DAGs for simplicity.
@@ 3.2 Language Multi-Agent RL
As shown in Figure 2, the essential stage of CollabUIAgents is MARL, after agentic fine-tuning the base model into a base UIAgent (Appendix C). We introduce a novel multi-agent Credit Re-Assignment (CR) strategy for better MARL. This approach utilizes the critic agent to provide fine-grained rewards at both the agent and conversation round levels, enabling agents to learn effectively from general knowledge of LLMs rather than the sparse environment-specfic reward, and improve collaboration by preference optimization with synthesized data from the adversarial agent. We also introduce an Edge Update trick to better accommodate trained LLMs in the multi-agent system.
Credit Re-Assignment  Traditional MARL systems use value decomposition for credit assignment (Ma & Luo, 2022), and recent work also uses multiple value estimators to alleviate reward sparcity in interactive environments (Bai et al., 2024). However, training critic models for each agent based on outcome rewards is computationally expensive and may hinder generalization. Instead, using an LLM-based critic agent, represented by $\pi_{\mathrm{critic}}$, we provide process rewards in a finer granularity for each agent in each conversation round. The action matrix $\bm{A}_{t}$ is assessed by the critic agent based on the observed environment state $o_{t}$, interaction history $H_{t-1}$, and the task query $q$, generating a reward matrix $\bm{R}_{t}=(r_{t}^{i,j})$, i=1,,n,j=1,,mi=1,...,n,j=1,...,m, where $r_{t}^{i,j}\in\{0,1\}$ denotes the intermediate reward for $a_{t}^{i,j}$:
@@
where $\pi_{\mathrm{critic}}(\cdot,\cdot,\cdot,\cdot)\in[0,1]^{n\times m}$. In our experiments, we use a strong gpt-4o-2024-08-06 to guide the CR process for better validity of the assigned rewards. However, the judgement accuracy would not be perfect when the complexity of the environment increases. To overcome the potential incorrect CR results, we introduce an adversarial agent $\pi_{\mathrm{adv}}$ to synthesize preference data for policy optimization substituting value-based learning. Specially, $\pi_{\mathrm{adv}}$ generates low-quality responses $a_{t}^{i,j,-}$ paired with $a_{t}^{i,j}$:
@@
For better clarity, the adversarial agent is only used for generating inferior actions and responses to help learning, instead of applying attacks (Pinto et al., 2017; Bukharin et al., 2023; Yuan et al., 2023; Lee et al., 2025). Here is the rationale of the design: (1) The critic agent provides a more detailed reward signal for each agent without training multiple value estimators; (2) The critic agent’s assessment is based on general pretrained LLMs, which by practice could expand the data scale by restoring intermediate decisions from failed trajectories according to the training environment, and thus may enhance performance; (3) Although errors in CR are inevitable, the synthesized preference data can still provide agents with meaningful preference signals. For example, even if $a_{t}^{i,j}$ is not the optimal action, it is still better than $a_{t}^{i,j,-}$. Indeed, environment-specific actions beyond general knowledge of the LLM might be misjudged. We argue that the situation is similar when a value estimator is poorly learned, where the policy might be optimized to a wrong direction, but our approach could keep rolling out new trajectories based on these actions without unlearning them. The qualitative study shown in Appendix B demonstrates CR results could be valid without the outcome reward.
MARL with Edge Updates  Different from classic MARL settings, agents in CollabUIAgents could communicate and the message network should also be rolled out in the optimization. To alleviate the overhead of learning the optimal network from enormous combinations of edges, we introduce an edge update trick, that randomly update edges to form a DAG message network, independently from model parameter updates. Through this process, we encourage agents to learn the awareness of multi-agent collaboration and adapt to diverse message networks rather than being rigid in locally optimal DAG pattern. As shown in Figure 2, the edge update is functioned before rolling out actions from the policy system to coordinate agents in a randomly sampled communication graph. Empirical evidence for its effectiveness is shown in Section 4.3. The overall learning objective for each agent $\pi_{i}$ is formulated as:
@@
where $\hat{H}_{t}^{i}=\{H_{t-1},C_{t}^{i,j},q\}$, $\theta_{i}$ are the updating parameters of agent $\pi_{i}$ and $\mathrm{ref}_{i}$ is the original model used as reference policy, $K_{|\mathcal{G}|}$ is a fully connected graph of $|\mathcal{G}|$ nodes, $\mathcal{E}_{\mathcal{G}}^{\prime}$ represents a DAG subgraph randomly sampled from $K_{|\mathcal{G}|}$, $\mathcal{P}(\mathcal{G},\mathcal{E}_{\mathcal{G}}^{\prime})$ is the preference dataset sampled with agents in the message network $\mathcal{E}_{\mathcal{G}}^{\prime}$, $\sigma$ is the sigmoid function, $\beta$ is the hyper-parameter, and $\pi_{\theta},\pi_{\mathrm{ref}}$ are the base model and reference model. This objective encourages the policy $\pi_{i}$ to assign higher probabilities to preferred actions $a_{t}^{i,j}$ compared to adversarial actions $a_{t}^{i,j,-}$. The policy is updated online and off-policy, similar to previous work (Bai et al., 2024).
@@ 3.3 Cross-Environment Adaptation
Direct Transfer  In scenarios where the new environment shares similarities with the training environment, agents can be deployed directly without additional training. For example, agents trained in mobile UI environments can directly apply their knowledge to web environments, leveraging the knowledge of common interaction patterns and UI elements. The multi-agent setup and according MARL stage are keys to decrease error rates through enhancing generalizable collaborative behaviors in agents as expected. The effectiveness is shown in Section 4.2 (applying CollabUIAgentsmobile to web environments).
[…]
Method #Params/#Agents Input SRAndroidWorld SRMMiniWoB++ $\boldsymbol{\Delta}$Generalization
@@
M3A (GPT-4) N/A Text 30.6 59.7 -
M3A (GPT-4) N/A Multimodal 25.4 67.7 -
SeeAct (GPT-4) N/A Multimodal 15.5 66.1 -
M3A (Gemini 1.5 Pro) N/A Text 19.4 57.4 -
M3A (Gemini 1.5 Pro) N/A Multimodal 22.8 40.3 -
@@
Qwen2 7B/1A Text 06.2 12.9 -
Qwen2 7B/4A Text 04.2 15.2 -
Qwen2 VL 2B/1A Multimodal 00.0 10.0 -
Qwen2 VL 2B/4A Multimodal 00.0 11.5 -
Qwen2.5 VL 3B/1A Multimodal 10.1 18.7 -
Qwen2.5 VL 3B/4A Multimodal 12.6 21.3 -
InfiGUIAgent (Qwen2 VL) 2B/1A Multimodal 09.1 15.6 05.6
DigiRL (Qwen2.5 VL) 3B/1A Multimodal 22.3 35.2 16.5
DigiRL (Qwen2.5 VL) 3B/4A Multimodal 20.1 38.7 20.0
@@
UIAgent (Qwen2) 7B/1A Text 18.9 48.4 35.5
UIAgent (Qwen2) 7B/4A Text 21.4 53.2 40.3
UIAgent (Qwen2) 7B/6A Text 18.9 54.3 41.5
CollabUIAgentsmobile (Qwen2) 7B/4A Text 29.3 61.2 48.3
CollabUIAgentsmobile (Qwen2) 7B/6A Text 32.7 59.7 46.9
Table 1: Experimental results on mobile operation environments. Success rates (SR) in AndoridWorld and MobileMiniWoB++ (MMiniWoB++) are listed. $\boldsymbol{\Delta}$Generalization indicates the performance gap between the base model and agent learning methods based on the model in MobileMiniWoB++. “7B/4A” denotes a four-agent system upon a 7B model.
@@ 4.1 Experimental Settings
Environments  We conduct experiments in both mobile operation and web browsing environments. For the mobile environments, we use AndroidWorld (Rawles et al., 2025) for training and MobileMiniWoB++ (Rawles et al., 2025) for testing: (1) AndroidWorld has 116 programmatic tasks across 20 real-world apps, such as Chrome, Markor, and Pro Expense. (2) MobileMiniWoB++ is derived from MiniWoB++ (Shi et al., 2017), which is a web-based benchmark. MobileMiniWoB++ shares the same observation space as AndroidWorld and supports 92 tasks from MiniWoB++. We use the success rate (SR) as an evaluation metric. For the web environments, we leverage Mind2Web (Deng et al., 2023) for training and AutoWebBench (Lai et al., 2024) for testing: (1) Mind2Web features over 2,000 open-ended tasks sourced from 137 websites in 31 different domains. (2) AutoWebBench is a bilingual benchmark featuring approximately 10,000 traces, from mainstream Chinese and English websites, providing a diverse dataset for web browsing. We use the step-success rate (SSR) as the evaluation metric. For agent learning methods on open-source LLMs, we use the performance gap between the base model and the trained model in unseen environments, $\boldsymbol{\Delta}$Generalization, to indicate generalizability for each agent learning method.
@@
4.2 Main Results
Method #Params/#Agents Input SSRMind2Web SSRAutoWebBench $\boldsymbol{\Delta}$Generalization